OpenStudy (anonymous):

Find the work done by a force of 4 pounds acting in the direction of 41° to the horizontal in moving an object 5 feet from (0, 0) to (5, 0). A. 13.1 ft-lb B. 16.2 ft-lb C. 15.1 ft-lb D. 30.2 ft-lb

OpenStudy (ddcamp):

 W = F*D = 4 * cos(41°) * 5 = 15.1 ft-lbs

OpenStudy (anonymous):

wow thank you so much you are a life saver ;-) @DDCamp